The length of a rectangle is 7 less than 3 times the width. If the perimeter is 34, find the dimensions.

Answer by checkley71(8403)

PERIMETER=2L+2W
L=3W-7
2(3W-7)+2W=34
6W-14+2W=34
8W=34+14
8W=48
W=48/8
W=6 ANSWER FOR THE WIDTH.
L=3*6-7
L=18-7
L=11 ANSWER FOR THE LENGTH.
PROOF
2*11+2*6=34
22+12=34
34=34